Give the total cost function



Tc(q)-Q2+10Q+100

1
Expert's answer
2022-01-31T09:44:24-0500

"TC = Q^2 + 10Q +100"

1) "TC = TVC + TFC;" "TVC = Q^2 +10Q;" "TFC = 100";

"AC = TC/Q = Q+10 +\frac{100}{Q}"

"AVC= TVC\/Q = Q+10"

2)"AC' = 1 - 100\times \frac{1}{Q^2} =0" => "\frac{1}{Q^2} = \frac{1}{100} => Q = 10"

"AC_{min} = 10 + 10 + \frac{100}{10}=30"